Headland Company sponsors a defined benefit pension plan for its employees. The following data relate to the operation of the plan for the years 2020 and 2021.

2020

2021

Projected benefit obligation, January 1 $604,400
Plan assets (fair value and market-related value), January 1 413,300
Pension asset/liability, January 1 191,100 Cr.
Prior service cost, January 1 159,200
Service cost 40,000 $58,500
Settlement rate 10 % 10 %
Expected rate of return 10 % 10 %
Actual return on plan assets 36,000 61,600
Amortization of prior service cost 69,300 50,000
Annual contributions 96,700 81,400
Benefits paid retirees 31,600 54,040
Increase in projected benefit obligation due to changes in actuarial assumptions 86,300 0
Accumulated benefit obligation at December 31 722,000 792,600
Average service life of all employees 20 years
Vested benefit obligation at December 31

465,900

1.Prepare a pension worksheet presenting both years 2020 and 2021.

2.

Prepare the journal entries (from the worksheet) to reflect all pension plan transactions and events at December 31 of each year.

3. For 2021, indicate the pension amounts reported in the financial statements.
Answer #1

Headland Company

Pension worksheet 2020-2021

General Entries

Memo Record

Items

Annual Pension expense

Cash

OCI- Prior service cost

OCI- Gain/Loss

Pension Asset/Liabilities

Project benefit obligation

Plan assets

Balance January 1 2020

191,100

Cr

604,400

Cr

413,300

Dr

Service cost

40,000

Dr

40,000

Cr

Interest cost

60,440

Dr

60,440

Cr

Actual Return

36,000

Cr

36,000

Dr

Unexpected loss

5,330

Cr

5,330

Dr

Amortization of PSC

69,300

Dr

69,300

Cr

Contributions

96,700

Cr

96,700

Dr

Benefits

31,600

Dr

31,600

Cr

Increase in PBO

86,300

Dr

86,300

Cr

Journal Entry for 2020

128,410

Dr

96,700

Cr

69,300

Cr

91,630

Dr

54,040

Cr

Accumulated OCI Dec 31, 2019

159,200

Dr

Balance Dec 31, 2020

89,900

Dr

91,630

Dr

245,140

Cr

759,540

Cr

514,400

Dr

Service cost

58,500

Dr

58,500

Cr

Interest cost

75,954

Dr

75,954

Cr

Actual return

61,600

Cr

61,600

Dr

Unexpected gain

10,160

Dr

10,160

Cr

Amortization of PSC

50,000

Dr

50000

Cr

Amortization of loss

784

Dr

784

Cr

Contributions

81,400

Cr

81,400

Dr

Benefits

54,040

Dr

54,040

Cr

Journal entry for 2021

133,798

Dr

81,400

Cr

50,000

Cr

10,944

Cr

8,546

Dr

Accumulated OCI Dec 31, 2020

89,900

Dr

91,630

Dr

Balance Dec 31, 2021

39,900

Dr

80,686

Dr

236,594

Cr

839,954

Cr

603,360

Dr

WORKING NOTES

1) Interest cost = (projected benefit obligation at the beginning of year * interest rate)

2017 = (604400 * 10%) = 60440

2018 = (759,540 * 10%) = 75,954

2) Calculation of unexpected loss or gain

Expected return = plan assets at the beginning of the year * Expected return

2017 = (413,300 * 10%)

41330

Actual return                  

(36,000)

Unexpected loss            

5330

2018 = (514,400 * 10%)

51,440

Actual return                  

(61,600)

Unexpected loss /gain         

(10,160)

3) Amortization of loss

2021 corridor test

Accumulated net gain/loss at the beginning of the year

91,630

10% of larger of PBO or plan asset i.e. higher of (759,540 * 10%) or (514,400 * 10%)

(75,954)

Amortizable Amount

15,676

2018 amortization (15,676 / 20years)

784

JOURNAL ENTRIES

Date

Particulars

Debit

Credit

Dec31,2020

Pension expense

128410

Other comprehensive income (G/L)

91630

              To cash

96700

              To other comprehensive income PSC  

69300

              To pension assets/ liabilities

54040

(to record pension related transactions)

Dec31,2021

Pension expense

133798

To pension assets/ liabilities

8546

              To cash

81400

               To other comprehensive income PSC  

50000

               Other comprehensive income (G/L)

10944

(to record pension related transactions)

Headland company

Income statement (partial)

For the year ended December 31, 2021

Expenses

Pension expense

133798

Headland company

Comprehensive Income statement (partial)

For the year ended December 31, 2021

Net income

Other comprehensive income (loss)

Asset gain (loss)

10160

Amortization of loss

784

Prior service cost amortization

50000

Comprehensive income

60,944

Headland company

Partial balance sheet

December 31, 2021

Liabilities

Pension liability

236,594

Stock holder’s equity

Accumulated other comprehensive loss (PSC)

39,900

Accumulated other comprehensive loss (G/L)

80,686